Colson Company has a line of credit with Federal Bank. Colson can borrow up to $434,500 at any time over the course of the 2018 calendar year. The following table shows the prime rate expressed as an annual percentage along with the amounts borrowed and repaid during the first four months of the year 2018. Colson agreed to pay interest at an annual rate equal to 3.00 percent above the bank's prime rate. Funds are borrowed or repaid on the first day of the month. Interest is payable in cash on the last day of the month. The interest rate is applied to the outstanding monthly balance. For example, Colson pays 7.25 percent (4.25 percent + 3.00 percent) annual interest on $79,100 for the month of January.

Month                                       Amount Borrowed or Repaid                 Prime Rate for the Month

January $79,100 4.25%

February 123,200 3.25%

March                                             (21,500) 3.75%

April                                                 28,000                                                       4.25%

a) Compute the amount of interest that Colson will pay on the line of credit for the first four months of 2018. (Round answers to nearest whole dollar)
Answer #1

Calculation of Interest amount

For the Month of January

Outstanding for the Month = $79100

Interest rate for the Month = 4.25 + 3 = 7.25%

Interest Amount for the Month = $79100 * 7.25% / 12

= $478

For the Month of February

Outstanding for the Month = $79100 + $123200 = 202300

Interest rate for the Month = 3.25 + 3 = 6.25%

Interest Amount for the Month = $202300 * 6.25% / 12

= $1054

For the Month of March

Outstanding for the Month = $79100 + $123200 - $21500 = 180800

Interest rate for the Month = 3.75 + 3 = 6.75%

Interest Amount for the Month = $180800 * 6.75% / 12

= $1017

For the Month of April

Outstanding for the Month = $79100 + $123200 - $21500 + $28000 = 208800

Interest rate for the Month = 4.25 + 3 = 7.25%

Interest Amount for the Month = $208800 * 7.25% / 12

= $1262

Interest is Paid in Cash.